### v1

# Drawing strategies in Strong Ramsey games for 3-uniform hypergraphs

Abstract.
The Strong Ramsey game $\mathcal{R}(B,G)$ is a two player game with players $P_{1}$ and $P_{2}$, where $B$ and $G$ are $k$-uniform hypergraphs for some $k\geq 2$. $G$ is always finite, while $B$ may be infinite. $P_{1}$ and $P_{2}$ alternately color uncolored edges $e\in B$ in their respective color and $P_{1}$ begins. Whoever completes a monochromatic copy of $G$ in their own color first, wins the game. If no one claims a monochromatic copy of $G$ in a finite number of moves, the game is declared a draw. In this paper, we give an infinite set of 3-uniform hypergraphs $\{G_{t}\}_{t\geq 3}$, such that $P_{2}$ has a drawing strategy in the Strong Ramsey game $\mathcal{R}(K_{\aleph_{0}}^{(3)},G_{t})$. This improves a result by David, Hartarsky and Tiba [undefe].
1. Introduction
One of the first people to study the Strong Ramsey game was Beck. In [undefb], he raised the question, whether $P_{1}$ has a winning strategy in $\mathcal{R}(K_{\aleph_{0}},K_{k})$ for all $k\geq 4$ and $\aleph_{0}$ being the smallest infinite cardinal. For $k=4$, his paper contains a winning strategy for $P_{1}$, which later turned out to be incomplete. Bowler and Gut [undefc] fixed the argument by using a computer algorithm. Note that $P_{2}$ does not have a winning strategy in $\mathcal{R}(B,G)$ for any graphs $B,G$, due to a simple strategy stealing trick (see [undeff, Theorem 1.3.1]).
In the meantime, Hefetz, Kusch, Narins, Pokrovskiy, Requilé and Sarid [undefg] constructed a $5$-uniform hypergraph $G^{*}$ for which they proved that $\mathcal{R}(K^{(5)}_{\aleph_{0}},G^{*})$ is a draw. Moreover, they suggested to tackle the following problem as an intermediate step towards Beck’s question:
Open problem 1.1.
Does there exist a graph $G$, such that $\mathcal{R}(K_{\aleph_{0}},G)$ is a draw?
For $l\geq 3$, let $\hat{K}_{2,l}$ be the graph obtained from $K_{2,l}$ together with the edge connecting the two vertices of degree $l$. For a graph $G$, let $G^{(k)}$ denote the $k$-uniform hypergraph, which can be obtained by adding the same fixed set of k2k-2 vertices to every edge in $E(G)$. In 2020, David, Hartarsky and Tiba [undefe] improved the above result by proving that $P_{2}$ has a drawing strategy in $\mathcal{R}(K_{\aleph_{0}}^{(4)},\hat{K}_{2,4}^{(4)})$. In Section 3 of this paper, we improve their result even further. Formally, we show that $P_{2}$ has a drawing strategy in $\mathcal{R}(K_{\aleph_{0}}^{(3)},\hat{K}_{2,4}^{(3)})$, thereby giving the first example of a 3-uniform hypergraph for which the corresponding Strong Ramsey game is a draw. This also supports their conjecture that $\mathcal{R}(K_{\aleph_{0}},\hat{K}_{2,4})$ is a draw. Moreover, we proved in [undefd] that, if $\mathcal{R}(K_{\aleph_{0}},\hat{K}_{2,4})$ were a draw, it would be a minimal example of this type.
For $s\geq 0$, $t\geq 3$, let $K_{2,t}(s)$ be the graph obtained by identifying the center of an $s$-star with one of the vertices of degree $t$ of $K_{2,t}$. Recently, Ai, Gao, Xu and Yan [undef] proved that $P_{2}$ has a drawing strategy in the Strong Ramsey game played on two disjoint boards $\mathcal{R}(K_{\aleph_{0}}\sqcup K_{\aleph_{0}},K_{2,t+1}(t-2))$ for all $t\geq 3$. Building on their work, we prove in Section 4 that $\mathcal{R}(K^{(3)}_{\aleph_{0}},K^{(3)}_{2,t+1}(t-2))$ is a draw for all $t\geq 3$. This suggests that the graphs $K_{2,t+1}(t-2)$ are promising candidates for a draw in the game $\mathcal{R}(K_{\aleph_{0}},K_{2,t+1}(t-2))$.
(a) The graph $\hat{K}_{2,4}$
(b) The graph $K_{2,3}(3)$
Figure 1.
2. Notation, Definitions and Basic Results
All (hyper)graphs considered in this paper are undirected and simple. Hence, we denote kk-(hyper-)edges as $x_{1}x_{2}\ldots x_{k}$ instead of $\{x_{1},x_{2},\ldots,x_{k}\}$ for clarity of the presentation. In the Strong Ramsey game $\mathcal{R}(B,G)$, we will call $B$ the board and $G$ the target graph. We refer to $P_{1}$ as she and $P_{2}$ as he and we will often say that a player claims, picks or takes an edge instead of colors it in their respective color. For brevity, we sometimes refer to an edge claimed by $P_{1}$ or $P_{2}$ as $P_{1}$-edge or $P_{2}$-edge, respectively. We use a similar notion, for graphs, which a player has claimed. We might also say that a player has or has built a graph $H$, meaning that the player has claimed all edges of a copy of $H\subseteq B$. Moreover, we say a vertex $x\in B$ is fresh, if no player has claimed an edge incident to $x$. For a given target graph $G$, we say that $P_{1}$ has a threat or $P_{1}$ threatens $P_{2}$, if $P_{1}$ has claimed a copy $H\subseteq B$ of GeG-e for an edge $e\in E(G)$, such that the corresponding $\hat{e}\in B$ is not yet claimed by either player, and similarly for $P_{2}$. We will call $H$ a threat graph in that case. In order to be concise, we will not give a formal definition of strategy here. Instead, we refer the reader to [undefa, Appendix C]. We define $k]\coloneqq\{1,\dots,k\}$.
We call the vertices added in order to obtain $G^{(k)}$ from $G$ the centers of $G^{(k)}$. On the other hand, in the case of $G=K_{\aleph_{0}}$ and $k=3$, given a vertex $x\in V(K_{\aleph_{0}}^{(3)})$, observe that the set $\{xyz\ |\ y,z\in V(K_{\aleph_{0}}^{(3)})\}$ spans $E(K_{\aleph_{0}})$. In this case, we will also refer to $x$ as center and we will call $\{yz\ |\ xyz\in E(K_{\aleph_{0}}^{(3)})\}$ the $x$-board. In both $K_{2,t}(s)$ and $\hat{K}_{2,l}$, all vertices of degree at least 3 will be called main vertices and all vertices of degree at most 2, minor vertices. Moreover in $\hat{K}_{2,l}$, the edge connecting the two vertices of degree $l+1$ is called the main edge.
In the course of this paper, it will be convenient to use time stamps in order to clarify which stage of the game we are talking about. Therefore, for a fixed point in time $T$, we define $E_{T}(P_{1})$ as the set of edges and $e_{T}(P_{1})$ as the number of edges that $P_{1}$ has claimed at $T$. For vertices $v,w\in B$, we let $d_{T}^{P_{1}}(v)$ be the number of edges from $E_{T}(P_{1})$, which are incident to $v$. Moreover, we define $\Delta_{T}^{P_{1}}(B)\coloneqq\max_{v\in B}\{d_{T}^{P_{1}}(v)\}$.
In the proof of Theorem 3.1 and Theorem 4.2, $P_{2}$ must react appropriately in order to prevent $P_{1}$ from claiming a copy of $G$. It will turn out to be helpful to distinguish cases based on how much progress $P_{1}$ has made toward claiming such a copy of $G$. For this, we will introduce the following progress bar.
Definition 2.1.
For a given target graph $G^{*}$ and board $B$ at a given point in time $T$, we define
$e_{T}^{P_{1}}(G^{*})=\max_{\mathclap{\begin{subarray}{c}H\subseteq B\\ H\cong G^{*}\\ E(H)\cap E_{T}(P_{2})=\emptyset\end{subarray}}}\{|E_{T}(P_{1})\cap E(H)|\}.$
From the point in time $T$ onward, notice that it takes $P_{1}$ at least $e(G^{*})-e_{T}^{P_{1}}(G^{*})$ moves to claim a copy of $G^{*}$. In particular, she has a threat at $T$ if and only if $e(G^{*})-e_{T}^{P_{1}}(G^{*})=1$. All the expressions above are defined similarly for $P_{2}$, and we might omit $T$ in case it is clear from the context.
In this paper, we are going to use a particular kind of figure to depict some board states in order to give the reader a visual intuition. In these figures, we will use the color violet only for edges of $P_{1}$ and blue only for edges of $P_{2}$. Moreover, all figures in this paper depict a board state after a move of $P_{1}$. Since drawing figures of 3-uniform hypergraphs is often a nightmare, we will only depict $x$-boards for specific $x\in B$, which we will indicate at the top of the figure. Hence, all the edges that are drawn as $2$-uniform edges correspond to 3-uniform edges containing the center of the corresponding board. In case, $P_{1}$ has claimed an edge $z_{1}z_{2}z_{3}$ not containing $x$, but the $x$-board is depicted in the figure, we will denote this by $z_{1}z_{2}z_{3}$ in the right corner of the figure. We will also use this notation, if $P_{1}$ has claimed an edge $e^{*}$ with special properties. Say that $P_{1}$ has claimed $k$ additional edges for which we do not have any further information. We will denote this by +k+k. Note that, we will not indicate, if edges of $P_{2}$ are missing in the figure. Oftentimes, we add a label $T$ to the top left corner of a figure in order to indicate which point in time it depicts.
For example in Figure 2, the $a$-board is depicted at $T_{1}$, $P_{1}$ has claimed the edges $z_{1}z_{2}z_{3}$, $e^{*}$ with special properties and four additional edges that we do not have any information about whereas $P_{2}$ has claimed $abv_{1},abv_{2},abv_{3},abv_{4},abv_{5}$ so far.
3. $\mathcal{R}(K_{\aleph_{0}}^{(3)},\hat{K}_{2,4}^{(3)})$ is a draw
In this section, all we do is proving that $\mathcal{R}(K_{\aleph_{0}}^{(3)},\hat{K}_{2,4}^{(3)})$ is a draw. For convenience, we let $G=\hat{K}_{2,4}^{(3)}$ in this section and we will refer to $G^{-}$ as the subgraph of $G$ with 8 edges, that is not the $K_{2,4}^{(3)}$.
Theorem 3.1.
$\mathcal{R}(K_{\aleph_{0}}^{(3)},G)$ is a draw.
In order to prove a draw in a Strong Ramsey game, from a certain board state onward, $P_{2}$’s strategy has to distract $P_{1}$ constantly from claiming a copy of $G$. This is only possible if $P_{2}$ threatens to claim a copy of $G$ himself. In the following lemma, we present a type of board state for which $P_{2}$ can make use of such a distraction strategy in order to force at least a draw. The idea of distracting $P_{1}$ was first used in [undefg] and it is essential for the proof of Theorem 3.1, since we only have to check the conditions required by this lemma, instead of playing the game for a countable number of moves, in order to guarantee a draw for $P_{2}$. Note that in the proof of Theorem 3.1, we always make use of the Distraction Lemma (3.2), whenever we state that $P_{2}$ has a drawing strategy from a given board state onward.
Lemma 3.2 (Distraction Lemma).
Suppose $P_{2}$ has claimed a $\hat{K}_{2,3}^{(3)}$ with main vertices $x,y$ and center $c$ in the game $\mathcal{R}(K_{\aleph_{0}}^{(3)},G)$, such that
  1. (i)
    $P_{1}$ does not have a threat;
  2. (ii)
    $P_{1}$ does not have a $\hat{K}_{2,3}^{(3)}$ with center $c$ and main vertex $x$;
  3. (iii)
    $P_{1}$ does not have a $\hat{K}_{2,3}^{(3)}$ with center $x$ and main vertex $c$.
If it is $P_{2}$’s turn, he has a drawing strategy.
Proof.
Let $a_{1}$ be a fresh vertex. $P_{2}$ claims $cya_{1}$ and $P_{1}$ must block $P_{2}$’s threat by taking $cxa_{1}$, because she had no threat. If claiming $cxa_{1}$ provides her with a threat graph $H$, we have $cxa_{1}\in E(H)$. Note that every edge of $H$ is incident to the center and one of the main vertices. Since $d^{P_{1}}(a_{1})=1$, it can neither be a center nor a main vertex of $H$. So $c$ and $x$ must be the center and one of the main vertices of $H$, but every threat graph with a vertex of degree 1 contains a $\hat{K}_{2,3}^{(3)}$, which was excluded by the assumptions $\mathit{(ii)}$ and $\mathit{(iii)}$.
$P_{2}$ continues by claiming $cya_{i}$ for fresh vertices $a_{i}$ as long as the game does not end or he can win by claiming $cxa_{i}$. Assume for a contradiction that $P_{1}$ won the game eventually. Then, she must have obtained a threat graph $H$ by blocking with $cxa_{i}$. Using the same argument as above, $P_{1}$ must have claimed a copy of $\hat{K}_{2,3}^{(3)}$ of type $\mathit{(ii)}$ or $\mathit{(iii)}$, which was forbidden. Hence, $P_{2}$ has a drawing strategy. ∎
Corollary 3.3.
In the game $\mathcal{R}(K_{\aleph_{0}}^{(3)},G)$, suppose $P_{2}$ has built a $\hat{K}_{2,3}^{(3)}$, such that $P_{1}$ has claimed at most 10 edges and she has no threat. If it is $P_{2}$’s turn, he has a drawing strategy.
Proof.
Let $x,y$ be the main vertices and $c$ the center of the $\hat{K}_{2,3}^{(3)}$, which $P_{2}$ has built and assume that either $\mathit{(ii)}$ or $\mathit{(iii)}$ of the Distraction Lemma (3.2) does not hold. In this case, $P_{1}$ has claimed at least seven edges, such that none of them is incident to $y$, since $cxy\in E(P_{2})$. $P_{1}$ can not have claimed another $\hat{K}_{2,3}^{(3)}$ with either main vertex or center $y$, since this requires at least four more edges on the $y$-board. Hence, $P_{1}$ has not claimed a copy of $\hat{K}_{2,3}^{(3)}$ with center $c$ or $y$ and main vertex $y$ or $c$, respectively. If we exchange $x$ and $y$ in the Distraction Lemma (3.2), we obtain that $P_{2}$ can force at least a draw. ∎
With those auxiliary results, we are finally able to prove the main theorem of this section.
Proof of Theorem 3.1.
$P_{1}$ claims an edge $z_{1}z_{2}z_{3}$. In his first five moves $P_{2}$ takes $abv_{i}$ for fresh vertices $a$, $b$, $v_{i}$ with $i\in[5]$. Let $e^{*}$ be the second edge of $P_{1}$ and let $T_{1}$ be the point in time after her sixth move. We obtain the board state in Figure 2.
Figure 2. The $a$-board at $T_{1}$. Note that the $b$-board looks almost identical: We only need to exchange $a$ and $b$ in the figure.
For a vertex $v\in V(B)\backslash\{v_{1},\dots,v_{5}\}$, define
$F_{v}\coloneqq\{v_{i}\in\{v_{1},\dots,v_{5}\}\ |\ \forall j\neq i\colon vv_{i}v_{j}\notin E_{T_{1}}(P_{1})\}$
as set of vertices, which are not incident to a violet edge in the $K_{5}$ induced by $\{v_{1},\dots,v_{5}\}$ on the $v$-board. At $T_{1}$, we are in one of the following two cases.
Case 1: $|F_{a}|\geq 2$ or $|F_{b}|\geq 2$.
Case 2: $P_{1}$ has claimed the edge set $\{z_{1}z_{2}z_{3},e^{*},av_{i_{1}}v_{i_{2}},av_{i_{3}}v_{i_{4}},bv_{j_{1}}v_{j_{2}},bv_{j_{3}}v_{j_{4}}\}$ for pairwise different $i_{1},i_{2},i_{3},i_{4}\in[5]$ and $j_{1},j_{2},j_{3},j_{4}\in[5]$.
We first show that this case distinction is complete. Suppose $|F_{a}|\leq 1$ and $|F_{b}|\leq 1$. By choice of the $v_{i}$, $e^{*}$ is incident to at most one $v_{i}$. Since $P_{1}$ has only claimed six edges at $T_{1}$, the four edges of $P_{1}$ not specified yet must be of the form $bv_{j_{1}}v_{j_{2}},bv_{j_{3}}v_{j_{4}}$ and $av_{i_{1}}v_{i_{2}},av_{i_{3}}v_{i_{4}}$ in order to fulfill $|F_{a}|\leq 1$ and $|F_{b}|\leq 1$. Hence, we are Case 2.
We start with proving that $P_{2}$ can force at least a draw in the first case.
Proof of Case 1.
In this case, $P_{2}$ will build a threat graph, whose center $c$ he will choose as follows. Let $c\in\{a,b\}$ be the vertex at $T_{1}$, such that $|F_{c}|\geq 2$. If both $|F_{a}|\geq 2$ and $|F_{b}|\geq 2$, we choose $c$, such that either $c\notin e^{*}$ or both $a,b\in e^{*}$. We let $x\in\{a,b\}$ be the other vertex, which will become a main vertex of the threat graph $P_{2}$ builds.
$P_{2}$ claims $cv_{i}v_{j}$, such that $i,j\in F_{c}$ and $d^{P_{1}}_{T_{1}}(v_{i}),d^{P_{1}}_{T_{1}}(v_{j})$ are minimal. Up to relabeling, suppose $P_{2}$ took $cv_{4}v_{5}$ in his sixth move. Moreover, we can assume up to relabeling that $cv_{1}v_{5},cv_{2}v_{5},cv_{3}v_{5}\notin E(P_{1})$ after the seventh move of $P_{1}$. In his seventh move, $P_{2}$ takes $cv_{k}v_{5}$ for $k\in[3]$, such that $d^{P_{1}}(v_{k})$ is currently maximal. Again by relabeling, suppose $P_{2}$ has taken $cv_{3}v_{5}$. Let $T_{2}$ be the point in time after the eighth move of $P_{1}$ (see Figure 3). Note that one of $cv_{1}v_{5}$ and $cv_{2}v_{5}$ is not claimed by $P_{1}$ at $T_{2}$.
Figure 3. The $c$-board at $T_{2}$.
We distinguish between the following four cases.
Case 1.1: $e_{T_{2}}^{P_{1}}(G)\leq 6$.
By symmetry, assume that $cv_{2}v_{5}$ is not taken by $P_{1}$, so $P_{2}$ claims $cv_{2}v_{5}$ and therefore builds a $\hat{K}_{2,3}^{(3)}$. After her next move, $P_{1}$ cannot have a threat by assumption and since $P_{1}$ has claimed 9 edges so far, $P_{2}$ can force at least a draw by Corollary 3.3.
Case 1.2: $e_{T_{2}}^{P_{1}}(G)=8$ and $P_{1}$ has built a copy of $K_{2,4}^{(3)}$.
For simplicity, let $H$ be the copy of $K_{2,4}^{(3)}$, which $P_{1}$ built. Since $z_{1}z_{2}z_{3}\in E(H)$ and $e_{T_{2}}^{P_{1}}(G)=8$, observe that both $cv_{1}v_{5}$ and $cv_{2}v_{5}$ are unclaimed at $T_{2}$. $P_{2}$ begins by blocking the threat of $P_{1}$. Immediately after his move, one can observe that $e^{P_{1}}(G)=4$ (by checking every combination of centers and main vertices). So as in Case 1.1, by symmetry $P_{2}$ picks $cv_{2}v_{5}$ in his next move, thereby claiming a $\hat{K}_{2,3}^{(3)}$. It is easy to see that $P_{1}$ cannot have a threat after her tenth move, since $e^{P_{1}}(G)=4$ two moves ago. By Corollary 3.3, $P_{2}$ obtains a drawing strategy.
Case 1.3: $e_{T_{2}}^{P_{1}}(G)=8$ and $P_{1}$ has built a copy of $G^{-}$.
Let $H$ be the copy of $G^{-}$ claimed by $P_{1}$. As in Case 1.2, we have $z_{1}z_{2}z_{3}\in E(H)$. So without loss of generality, let $z_{1}$ be the center of $H$, let $z_{2}$ be a main vertex and let $z^{*}$ be the other main vertex. Both $cv_{1}v_{5}$ and $cv_{2}v_{5}$ are not taken by either player.
$P_{2}$ blocks the threat of $P_{1}$. As long as $P_{1}$ creates threat graphs by claiming either $z_{1}z_{2}w_{i}$ or $z_{1}z^{*}w_{i}$ for vertices $w_{i}$, $P_{2}$ blocks by taking $z_{1}z^{*}w_{i}$ or $z_{1}z_{2}w_{i}$, respectively. Since the game ends in a draw, if she does not stop claiming edges of those type, eventually $P_{1}$ must claim an edge $\hat{e}$, which does not contain both $z_{1}$ and $z_{2}$ or both $z_{1}$ and $z^{*}$. Let $T_{3}$ be the point in time after $P_{1}$ has claimed $\hat{e}$. We obtain the board state in Figure 4 (from a $P_{1}$ perspective), where the left blob represents the vertices $v$ for which $P_{1}$ picked $z_{1}z_{2}v$ and $P_{2}$ blocked accordingly and the right blob represents all $w$ for which $P_{1}$ claimed $z_{1}z^{*}w$ and $P_{2}$ blocked accordingly.
Refer to caption Figure 4. The $z_{1}$-board with all edges of $P_{1}$ after she took $\hat{e}$. We will now prove that $e_{T_{3}}^{P_{1}}(G)=7$ and hence she has no threat. Assume for contradiction that after $P_{1}$ took $\hat{e}$, she has claimed a threat graph $H$. Then, $\hat{e}\in E(H)$. Additionally, the center and the main vertices of $H$ must have degree at least 4 and are therefore $z_{1},z_{2}$ and $z^{*}$. If $z_{1}$ was the center, then $z_{2}$ and $z^{*}$ have to be the main vertices of $H$. In this case, $\hat{e}$ is incident to either both $z_{1}$ and $z_{2}$ or both $z_{1}$ and $z^{*}$, but those possibilities were excluded. $z_{2}$ cannot be the center, because the $z_{2}$-board contains at most a $P_{1}$-star centered at $z_{1}$ and potentially $\hat{e}$. Hence, there cannot be a threat with center $z_{2}$, and similarly for $z^{*}$.
$P_{2}$ claims $cv_{2}v_{5}$ up to relabeling vertices. This yields him a $\hat{K}_{2,3}^{(3)}$. We have either $\hat{e}=cv_{1}v_{5}$ or $\hat{e}\neq cv_{1}v_{5}$. In the latter case, $P_{1}$ must have taken $cv_{1}v_{5}$ in her next move after $T_{3}$, because he wins on the next move otherwise.
If $\hat{e}\neq cv_{1}v_{5}$, let $T_{4}$ be the point in time after $P_{1}$ took $cv_{1}v_{5}$. We check the conditions of the Distraction Lemma (3.2) at $T_{4}$. Suppose $P_{1}$ has claimed a copy of a threat graph $H$. Since she had no threat a move ago (at $T_{3}$), we must have $cv_{1}v_{5}\in E(H)$. Since the only possible centers for $H$ are $z_{1},z_{2}$ and $z^{*}$, by choice of $c,v_{1},v_{5}$, $z^{*}$ must be the center of $H$. This is not possible as the $z^{*}$-board contains at most a $P_{1}$-star centered at $z_{1}$ and two additional edges of $P_{1}$. For conditions $\mathit{(ii)}$ and $\mathit{(iii)}$, observe that if $z^{*}\neq x$, then $x$ is contained in at most three edges of $P_{1}$. Hence, it can neither be a main vertex nor a center of a $\hat{K}_{2,3}^{(3)}$. If $z^{*}=x$, then $c$ and $v_{5}$ satisfy conditions $\mathit{(ii)}$ and $\mathit{(iii)}$ as the only possible centers for a $\hat{K}_{2,3}^{(3)}$ claimed by $P_{1}$ are $z_{1},z_{2}$ and $z^{*}$.
If $\hat{e}=cv_{1}v_{5}$, $P_{1}$ can begin a second threat phase, if she picks $z_{1}z_{2}w$ or $z_{1}z^{*}w$ for a vertex $w$. $P_{2}$ blocks accordingly as before until $P_{2}$ selects an edge $e^{*}$, which does not contain both $z_{1}$ and $z_{2}$ or both $z_{1}$ and $z^{*}$. This results in the board state from the last paragraph at $T_{4}$ up to the number of times, that $P_{1}$ threatened $P_{2}$ by taking $z_{1}z_{2}w$ or $z_{1}z^{*}w$, respectively. Since $P_{2}$ had a drawing strategy in this position, this completes Case 1.3.
Case 1.4: $e_{T_{2}}^{P_{1}}(G)=7$.
$P_{2}$ claims $cv_{2}v_{5}$ up to relabeling. Since $P_{1}$ had no threat, we know that $cv_{1}v_{5}\in E(P_{1})$ after her ninth move. If $e^{P_{1}}(G)\leq 7$, we get by Corollary 3.3 that $P_{2}$ can force at least a draw. So suppose that $e^{P_{1}}(G)=8$ after her ninth move and let $H$ be the threat graph of $P_{1}$. First of all, $P_{2}$ blocks the threat. Since $\{z_{1},z_{2},z_{3}\}\cap\{c,v_{1},v_{5}\}=\emptyset$, only one of those edges is contained in $E(H)$.
If $z_{1}z_{2}z_{3}\in E(H)$, we obtain a board state, which can be reached via Case 1.2 or Case 1.3 as follows. Suppose $H$ was a copy of $G^{-}$, then $P_{1}$ could have reached the same position by claiming $G^{-}$ in his first eight moves. After $P_{2}$ blocks her threat as in Case 1.3, $P_{1}$ takes $cv_{1}v_{5}$ and $P_{2}$ takes $cv_{2}v_{5}$. Similarly for Case 1.2, if $P_{1}$ built a copy of $K_{2,4}^{(3)}$ within her first eight moves instead.
Assume for contradiction that $cv_{1}v_{5}\in E(H)$. If $c$ was the center of $H$, then $d_{T_{1}}^{P_{1}}(c)=5$. If $|F_{x}|\leq 1$, this is not possible as $P_{1}$ has taken $z_{1}z_{2}z_{3}$ and $xv_{i}v_{j}$, which are both not on the $c$-board. Otherwise, $|F_{x}|\geq 2$ and recall that at the beginning of this proof, we have ensured that $e^{*}=cxw$ for a vertex $w$ in this case. We obtain $e^{*}\notin E(H)$, since either $v_{1}$ or $v_{5}$ has to be a main vertex of $H$, but both $cxv_{1}$ and $cxv_{5}$ are taken by $P_{2}$. If $v_{5}$ was the center of $H$, then $d_{T_{1}}^{P_{1}}(v_{5})=5$ by construction. In case $|F_{c}|<5$, we had $cv_{k_{1}}v_{k_{2}}\in E_{T_{1}}(P_{1})$ with $k_{1}\neq 5\neq k_{2}$ and therefore $d_{T_{1}}^{P_{1}}(v_{5})<5$. If otherwise $|F_{c}|=5$ and $d_{T_{1}}^{P_{1}}(v_{5})=5$, by the pigeonhole principle there exist $i,j\in[4]$, such that $d_{T_{1}}^{P_{1}}(v_{i}),d_{T_{1}}^{P_{1}}(v_{j})\leq 3$ contradicting the choice of $cv_{4}v_{5}$ and hence $v_{5}$ cannot be a center of $H$. If $v_{1}$ was the center of $H$, after $P_{1}$’s seventh move, we must have had $d^{P_{1}}(v_{1})=6$. Since every vertex of $G$ other than the center has degree less than five, this implies $d^{P_{1}}(v_{2}),d^{P_{1}}(v_{3})\leq 5$, which contradicts the choice of $cv_{3}v_{5}$.
This completes the proof of Case 1. ∎
It will be easier to show that $P_{2}$ can achieve at least a draw in the second case as it is more specific.
Proof of Case 2.
Suppose, we are at $T_{1}$ and $P_{1}$ has claimed the edges listed above. It is easy to check that $e_{T_{1}}^{P_{1}}(G)\leq 3$. Up to relabeling vertices, we can assume that $P_{1}$ has claimed $av_{1}v_{2}$ and $av_{3}v_{4}$ as in Figure 5.
Figure 5. The $a$-board at $T_{1}$ in Case 2.
Without loss of generality, $P_{2}$ claims $av_{4}v_{5}$ and $av_{1}v_{5}$ in his next two moves. Thereafter, $P_{2}$ claims either $av_{2}v_{5}$ or $av_{3}v_{5}$, if possible (see Figure 6(a)) or both $av_{2}v_{5}$ and $av_{3}v_{5}$ were claimed by $P_{1}$. In the latter case, $P_{2}$ claims $av_{1}v_{4}$ and then $P_{1}$ did not claim both $av_{1}v_{3}$ and $av_{2}v_{4}$ after her next move. Up to symmetry, $P_{2}$ claims $av_{2}v_{4}$ in his next move to obtain a $\hat{K}_{2,3}^{(3)}$. This yields the board state in Figure 6(b).
(a) $P_{1}$ must have taken $av_{3}v_{5}$, because $P_{2}$ wins on the next move otherwise.
Figure 6.
In Figure 6(a), one can check that for all $x\in V(B)$, the $x$-board contains at most six edges of $P_{1}$ and therefore, we have $e^{P_{1}}(G)\leq 6$, so $P_{2}$ has a drawing strategy by Corollary 3.3. From comparing the edge sets $P_{1}$ has claimed in Figure 6(a) and Figure 6(b), one can easily deduce that $e^{P_{1}}(G)\leq 7$ in Figure 6(b). Hence by Corollary 3.3, $P_{2}$ has a drawing strategy in Case 2. ∎
Since $P_{2}$ has a drawing strategy in Case 1 and Case 2 and those were all the cases that could occur, this completes the proof. ∎
4. $\mathcal{R}(K_{\aleph_{0}}^{(3)},K^{(3)}_{2,t+1}(t-2))$ is a draw for all $t\geq 3$
In this section, we are going to refine the main proof from Ai, Gao, Xu and Yan [undef] in order to show that $\mathcal{R}(K^{(3)}_{\aleph_{0}},K^{(3)}_{2,t+1}(t-2))$ is a draw for all $t\geq 3$. For convenience, we let $G_{t}=K^{(3)}_{2,t+1}(t-2)$ for the rest of this section. We begin by stating a lemma equivalent to the Distraction Lemma from Section 3. Due to their similarity, we will not prove it.
Lemma 4.1 (Distraction Lemma).
For $t\geq 3$, suppose $P_{2}$ has built a $K^{(3)}_{2,t}(t-2)$ with main vertices $x,y$ and center $c$ in the game $\mathcal{R}(K_{\aleph_{0}}^{(3)},G_{t})$, such that
  1. (i)
    $P_{1}$ does not have a threat;
  2. (ii)
    $P_{1}$ has not claimed a $K_{2,t}^{(3)}$ with center $c$ and main vertex $x$;
  3. (iii)
    $P_{1}$ has not claimed a $K_{2,t}^{(3)}$ with center $x$ and main vertex $c$.
If it is $P_{2}$’s turn, he has a drawing strategy. ∎
We continue with proving the main theorem of this section.
Theorem 4.2.
$\mathcal{R}(K_{\aleph_{0}}^{(3)},G_{t})$ is a draw for all $t\geq 3$.
Proof.
Fix $t\geq 3$ and let $B$ be the infinite complete 3-uniform board $K_{\aleph_{0}}^{(3)}$. Let $a_{1}a_{2}a_{3}$ be the first edge claimed by $P_{1}$. In his first 2t12t-1 moves, $P_{2}$ selects edges of the form $x_{1}x_{2}y_{i}$ for $i\in[2t-1]$, such that $x_{1}$, $x_{2}$ and all $y_{i}$ are chosen fresh at that time. Let $T_{1}$ be the point in time after $P_{1}$ has taken $2t$ edges.
If $e^{P_{1}}_{T_{1}}(G_{t})=2t$, fix a copy $H\subseteq B$ of $G_{t}$ with center and main vertices $c,m_{1},m_{2}\in B$, such that $E_{T_{1}}(P_{1})\subseteq E(H)$. Observe that the center and the main vertices of $H$ are uniquely determined, i.e. there exists no other copy at $T_{1}$ with a center other than $c$ and main vertices other than $m_{1}$ and $m_{2}$, such that $E_{T_{1}}(P_{1})\subseteq E(H)$. Moreover, we can assume without loss of generality, that $c=a_{1}$ and $m_{1}=a_{2}$ in that case.
We choose $x_{C}\in\{x_{1},x_{2}\}$ as follows. If $e^{P_{1}}_{T_{1}}(G_{t})=2t$ and $m_{2}=x_{1}$ or $m_{2}=x_{2}$, we choose $x_{C}=x_{2}$ or $x_{C}=x_{1}$, respectively and therefore guarantee, that $x_{C}$ is not a main vertex of $H$. In all other cases, we choose $x_{C}\in\{x_{1},x_{2}\}$, such that $d^{P_{1}}_{T_{1}}(x_{C})$ is minimal. $x_{C}$ will be the center of the $K^{(3)}_{2,t}(t-2)$, which $P_{2}$ will build in order to apply the Distraction Lemma (4.1). We will denote the other vertex in $\{x_{1},x_{2}\}$ with $x_{M}$ and it will become a main vertex of the $K^{(3)}_{2,t}(t-2)$. Note that, the only case in which we do not choose $d^{P_{1}}_{T_{1}}(x_{C})$ to be minimal is, if $P_{1}$ has claimed a (2t1)(2t-1)-star on the $a_{1}$-board with $x_{C}$ as a leaf together with $a_{1}x_{C}x_{M}$. $P_{2}$ now picks $x_{C}zy_{1}$ for a fresh vertex $z$. Let $T_{2}$ be the point in time after $P_{1}$ has made $2t+1$ moves.
(a) The $x_{C}$-board after $2t$ moves of $P_{1}$.
(b) The $x_{C}$-board after $2t+1$ moves of $P_{1}$.
Figure 7.
We obtain the following two cases.
Case 1: $e^{P_{1}}_{T_{2}}(G_{t})<2t+1$
Case 2: $e^{P_{1}}_{T_{2}}(G_{t})=2t+1$
We will show that in either case, $P_{2}$ has a drawing strategy.
Proof of Case 1.
In his next t1t-1 moves, $P_{2}$ selects t1t-1 edges of the form $x_{C}zy_{i}$ for $i\in[2t-1]$, which is possible due to the pigeonhole principle. Up to relabeling, say $P_{2}$ has taken $x_{C}zy_{1},\dots,x_{C}zy_{t}$ and let $T_{3}$ be the point in time after $P_{1}$ has made $3t$ moves. If $P_{2}$ cannot win immediately after $3t$ moves, $P_{1}$ must have claimed $x_{C}zy_{t+1},\dots,x_{C}zy_{2t-1}$, since $e^{P_{1}}_{T_{3}}(G_{t})<3t$. We verify that the conditions of the Distraction Lemma (4.1) hold.
Figure 8. The $x_{C}$-board after $3t$ moves by $P_{1}$.
If $P_{1}$ had a threat graph $H$ at $T_{3}$, then t1t-1 of the $t$ edges specified in Figure 8 must be contained in $H$. Since $\{a_{1},a_{2},a_{3}\}\cap\{x_{C},z,y_{i}\}=\emptyset$ for all $i\in\{t+1,\dots,2t-1\}$, in fact all such edges $x_{C}zy_{i}$ must be contained in $H$. By choice of $z$, $x_{C}$ must be the center of $H$. Therefore, $d_{T_{3}}^{P_{1}}(x_{C})=3t-1$ (it cannot be larger, because $x_{C}\notin\{a_{1},a_{2},a_{3}\}$) and $d_{T_{1}}^{P_{1}}(x_{C})=2t-1$. By choice of $x_{C}$, we have $d_{T_{1}}^{P_{1}}(x_{M})=2t-1$, so $P_{1}$ must have claimed 2t12t-1 edges of the form $x_{C}x_{M}v$ for $v\notin\{y_{1},\dots,y_{2t-1}\}$ at $T_{1}$. Since $|V(G_{t})|=2t+2$, but the edges $P_{1}$ has claimed cover at least $3t+1$ vertices of $B$, it is clear that there cannot be a threat. In order to verify $\mathit{(ii)}$ and $\mathit{(iii)}$ of the Distraction Lemma (4.1), observe that $z$ cannot be a center of a $K_{2,t}^{(3)}$ (with main vertex $x_{C}$) claimed by $P_{1}$, because it was chosen fresh in move $2t$ by $P_{2}$.
Suppose that $P_{1}$ has claimed a copy $H$ of $K_{2,t}^{(3)}$ at $T_{3}$ with center $x_{C}$ and main vertices $z,z_{M}$ for some vertex $z_{M}\in V(B)$. We have $x_{C}zy_{t+1},\dots,x_{C}zy_{2t-1}\in E(H)$, since $d_{T_{3}}^{P_{1}}(z)\leq t$ and hence $x_{M}\neq z_{M}$. Let $y^{*}$ be the other minor vertex of $H$. By choice of $z$, we know that in her last $t$ moves, $P_{1}$ must have picked $x_{C}zy_{t+1},\dots,x_{C}zy_{2t-1},x_{C}zy^{*}$. Therefore, the edges $x_{C}z_{M}y_{t+1},\dots,x_{C}z_{M}y_{2t-1},x_{C}z_{M}y^{*}$ had already been claimed by $P_{1}$ at $T_{1}$, so $d^{P_{1}}_{T_{1}}(x_{C})\geq t$. By choice of $x_{C}$, we have $d^{P_{1}}_{T_{1}}(x_{M})\geq t$. This implies $y^{*}=x_{M}$ and the t1t-1 edges not specified at $T_{1}$ must contain $x_{M}$, but cannot contain $x_{C}$. Then, there can be at most two edges at $T_{3}$ (namely $x_{C}z_{M}y^{*}$ and $x_{C}zy^{*}$), which are incident to both $x_{C}$ and $x_{M}$. Applying the Distraction Lemma (4.1) with $x_{C}$ and $x_{M}$ as main vertex and center provides us a drawing strategy for $P_{2}$. ∎
Proof of Case 2.
Let $H$ be a copy of a subgraph of $G_{t}$ with $2t+1$ edges, center $a_{1}$ and main vertices $a_{2},a_{M}$, which $P_{1}$ has claimed at $T_{2}$. For convenience, let
$E(H)$ $=\{a_{1}a_{2}v_{i},a_{1}a_{M}v_{i}\ |\ i\in\{1,\dots,s_{1}\}\}$
$\cup\{a_{1}a_{2}v_{i}\ |\ i\in\{s_{1}+1,\dots,s_{1}+s^{\prime}\}\}$
$\cup\{a_{1}a_{M}v_{i}\ |\ i\in\{s_{1}+s^{\prime}+1,\dots,s_{1}+s^{\prime}+s^{\prime\prime}\}\}.$
Define $s_{2}=s^{\prime}+s^{\prime\prime}$. Since $2s_{1}+s_{2}=2t+1$, we have that $s_{2}$ is odd and $s_{1}\leq t$. Let
$A=\{a_{1}a_{M}v_{i}\ |\ i\in\{s_{1}+1,\dots,s_{1}+s^{\prime}\}\}\cup\{a_{1}a_{2}v_{i}\ |\ i\in\{s_{1}+s^{\prime}+1,\dots,s_{1}+s^{\prime}+s^{\prime\prime}\}\}.$
Figure 9. The $a_{1}$-board from the perspective of $P_{1}$ at $T_{2}$. The left blob contains the vertices $v_{1},\dots,v_{s}$, the middle contains $v_{s+1},\dots,v_{s+s^{\prime}}$ and the right contains $v_{s+s^{\prime}+1},\dots,v_{s+s^{\prime}+s^{\prime\prime}}$. The dashed edges are the ones in $A$, which are claimed by neither player yet.
In his $(2t+1)^{\text{st}}$ move, $P_{2}$ picks any edge from $A$. As long as he can, $P_{2}$ plays according to the following algorithm. We will prove later that the first time $P_{2}$ cannot play according to the algorithm, the Distraction Lemma (4.1) is applicable and he can force at least a draw.
  1. (i)
    Whenever $P_{1}$ picks an unclaimed edge of $A$, $P_{2}$ picks any other unclaimed edge of $A$. (This is always possible, since $A$ is even after the $(2t+1)^{\text{st}}$ move of $P_{2}$.)
  2. (ii)
    If $P_{1}$ claims $a_{1}a_{2}v$ or $a_{1}a_{M}v$ for $v\notin\{a_{2},a_{M}\}$ and $a_{1}a_{2}v,a_{1}a_{M}v\notin A$, $P_{2}$ chooses $a_{1}a_{M}v$ or $a_{1}a_{2}v$ respectively.
  3. (iii)
    Otherwise, $P_{2}$ picks an unclaimed edge $x_{C}zy_{i}$ for $i\in[2t-1]$ as long as possible.
Assume the game is played up to $T_{4}$, the point in time when $P_{1}$ has either won or it is $P_{2}$’s move and he cannot play according to the algorithm. Let $E_{1}$ be the set of $P_{1}$-edges for which $P_{2}$ has claimed an edge on her next turn due to part $\mathit{(i)}$ of the algorithm and similarly for $E_{2},E_{3}$. Assume for a contradiction that $P_{1}$ won by building a copy $H$ of $G_{t}$ with center $c^{*}$ and main vertices $m_{1}^{*},m_{2}^{*}$, such that $d^{P_{1}}_{T_{4}}(m_{1}^{*})\geq 2t-1$. We begin by proving the following claim.
Claim: Either the center of $H$ is not $a_{1}$ or one of its main vertices is not $a_{2}$ or $a_{M}$.
Assume the contrary. By the calculations above, we obtain $|A|=2(t-s_{1})+1$. Due to part $\mathit{(i)}$ of the algorithm, $P_{1}$ can connect at most $t-s_{1}$ vertices from $v_{s_{1}+1},\dots,v_{s_{1}+s_{2}}$ to both $a_{2}$ and $a_{M}$. By $\mathit{(ii)}$, she can connect any other vertex to at most one of $a_{2},a_{M}$ on the $a_{1}$-board. So $P_{1}$ can obtain at most a $K_{2,t}$ and hence cannot even claim a $K_{2,t+1}^{(3)}$ with center $a_{1}$ and main vertices $a_{2}$ and $a_{M}$.
Note that at $T_{4}$ we have
$E_{1}\cup E_{2}\subseteq\{a_{1}a_{2}v^{\prime}\ |\ v^{\prime}\in V(B)\backslash\{a_{M}\}\}\cup\{a_{1}a_{M}v^{\prime\prime}\ |\ v^{\prime\prime}\in V(B)\backslash\{a_{2}\}\}.$
In addition, $|E_{3}|\leq t$, because $P_{2}$ has won by playing according to the algorithm otherwise.
Which vertex can $c^{*}$ possibly be? We know that $d_{T_{4}}^{P_{1}}(c^{*})\geq 3t$, hence the options for $c^{*}$ are $a_{1},a_{2}$ and $a_{M}$. Both the $a_{2}$-board and the $a_{M}$-board contain an arbitrarily large star centered at $a_{1}$ and potentially edges from $E_{3}$. But since $|E_{3}|\leq t$, this is not enough to build a copy of $G$, so $c^{*}=a_{1}$. Again by a degree argument, we have $m_{1}^{*}=a_{2}$ without loss of generality. ($m_{1}^{*}=a_{M}$ is almost the same.) Since $m_{2}^{*}\neq a_{M}$ by the claim above, $|\{e\in E_{1}\cup E_{2}\ |\ m_{2}^{*}\in e\}|\leq 2$ and $d_{T_{4}}^{P_{1}}(m_{2}^{*})\geq t+1$, we must have
(1) $\big|\{e\in E_{T_{4}}(P_{1})\cap E(H)\ |\ m_{2}^{*}\in e\}\backslash E_{3}\big|\in\{1,2\},$
so one of $a_{1}a_{2}m_{2}^{*},a_{1}a_{M}m_{2}^{*}$ must be contained in $E(H)$. Since $a_{1}a_{2}m_{2}^{*}\notin E_{T_{4}}(P_{1})\cap E(H)$, we obtain $a_{1}a_{M}m_{2}^{*}\in E_{T_{4}}(P_{1})\cap E(H)$. Moreover, all edges of $E_{3}$ must contain $a_{1}$ and $m_{2}^{*}$, because $|\{e\in E_{T_{4}}(P_{1})\cap E(H)\ |\ a_{1},m_{2}^{*}\in e\}|\leq t$, otherwise. In this case, $a_{1}a_{M}m_{2}^{*}\notin E(H)$, since $a_{1}a_{2}a_{M}\notin E_{T_{4}}(P_{1})$ contradicting Equation 1 and therefore our assumption.
So it must be $P_{2}$’s move and he cannot play according to the algorithm, due to $\mathit{(iii)}$. By the pigeonhole principle, $P_{2}$ has claimed the edges $x_{C}zy_{2},\dots,x_{C}zy_{t}$ without loss of generality and if he can not win at $T_{4}$, then $P_{1}$ must have claimed $x_{C}zy_{t+1},\dots,x_{C}zy_{2t-1}$ in the meantime. Since $P_{2}$ has claimed a $K^{(3)}_{2,t}(t-2)$ and it is his move, we check the conditions of the Distraction Lemma (4.1).
Figure 10. The actual board state at $T_{4}$ with the $a_{1}$-board on the left and the $x_{C}$-board on the right. Moreover, $P_{1}$ has claimed at most $(t-s_{1})$ edges of AA (the dashed edges as in Figure 9). Since $|E_{3}|=t$, note that all $P_{1}$-edges except one are displayed either in the left part or the right part of the figure.
Since $e^{P_{1}}_{T_{2}}(G_{t})=2t+1$, we had $e^{P_{1}}_{T_{1}}(G_{t})=2t$. Moreover, $a_{M}$ has been chosen by $P_{1}$ at latest in her $(2t)^{\textit{th}}$ move, so by choice of $x_{C}$, we have $x_{C}\notin\{a_{1},a_{2},a_{M}\}$. Since $z$ was chosen in move $2t$, we also have $z\notin\{a_{1},a_{2},a_{M}\}$. So $d^{P_{1}}_{T_{4}}(x_{C}),d^{P_{1}}_{T_{4}}(z)\leq t+2$ and since $t\geq 3$, there cannot be a $K_{2,t}^{(3)}$ claimed by $P_{1}$, with center $x_{C}$ or $z$ and hence $\mathit{(ii)}$ and $\mathit{(iii)}$ of the Distraction Lemma (4.1) hold.
If $P_{1}$ had a threat graph $H$ at $T_{4}$, then its center must be contained in at least 3t13t-1 edges. By the same degree argument as above, the only possible choices for its center are $a_{1},a_{2},a_{M}$. The $a_{2}$-board contains a $P_{1}$-star centered at $a_{1}$ and at most one additional $P_{1}$-edge, since $x_{C}zy_{t+1},\dots,x_{C}zy_{2t-1}\in E_{3}$. Since it could be that $a_{M}=y_{i}$ for some $y_{i}$, the $a_{M}$-board contains at most a $P_{1}$-star and two additional $P_{1}$-edges, so both $a_{2}$ and $a_{M}$ cannot be the centers of $H$. On the $a_{1}$-board, only $a_{2}$ and $a_{M}$ have degree at least 4 at $T_{4}$, so if $t\geq 4$, only $a_{2}$ and $a_{M}$ could be the main vertices of $H$. By definition of $T_{4}$ the last edge $P_{1}$ claimed before $T_{4}$ must be contained in $H$, but cannot be of the form $a_{1}a_{2}v$ or $a_{1}a_{M}v$ for some $v$, because in that case $P_{2}$ reacted with part $\mathit{(i)}$ or $\mathit{(ii)}$ of the algorithm. Therefore, $P_{1}$ cannot have a threat graph at $T_{4}$, if $t\geq 4$.
If $t=3$, the only type of vertex $v\notin\{a_{2},a_{M}\}$, which satisfies $d^{P_{1}}_{T_{4}}(v)=3$ on the $a_{1}$-board is connected with $a_{2}$ and $a_{M}$ and contained in the non-specified edge from $E_{3}$. Since either $a_{2}$ or $a_{M}$ must be the other main vertex of $H$, $a_{1}a_{2}v\notin E(H)$ or $a_{1}a_{M}v\notin E(H)$ respectively. Therefore, $|E(H)\cap\{e\in E_{T_{4}}(P_{1})\ |\ v\in e\}|\leq 2$ contradicting that $v$ was a main vertex of $H$. ∎
This shows that $P_{2}$ has also a drawing strategy in Case 2, which completes the proof. ∎